OpenStudy (anonymous):

find the slope of the line (2,1)(6,10)

OpenStudy (anonymous):

\[m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

OpenStudy (anonymous):

4/9

OpenStudy (anonymous):

\[\frac{ 10-1 }{ 6-2 }=\frac{ 9 }{ 4 }\]

OpenStudy (anonymous):

You used x where you should have used y and vice versa

OpenStudy (anonymous):

9/4